OpenStudy (richyw):

can somone help me with this derivative? \[\frac{d}{dz} (1+i)^{z^2}\] assume an appropriate branch cut when necessary.

zepdrix (zepdrix):

Ughhh I've got a test on this stuff next week :( I Better get studying!!

OpenStudy (richyw):

mine is tomorrow!

OpenStudy (anonymous):

you can try this with a u substitution if you know how? does that sound a little familiar?

zepdrix (zepdrix):

So we would use the differentiation rules for exponential functions yes? \[\Large\bf\sf =\quad (1+\mathcal i)^{z^2}\cdot \ln(1+\mathcal i)\frac{d}{dz}z^2\]This approach maybe? :O

zepdrix (zepdrix):

Woops I should call that \(\Large\bf\sf log(1+\mathcal i)\) And then we expand out the log after that. Mmmmm am I doing this right? +_+ Hopefully you have some idea...

OpenStudy (richyw):

I don't understand how you are getting that log?

zepdrix (zepdrix):

\[\Large\bf\sf (a^x)'\quad=\quad a^x \ln a\]

OpenStudy (anonymous):

If y = x^x then ln y = ln (x^x) so... lny = xlnx

zepdrix (zepdrix):

Remember back to taking the derivative of exponential functions? :) The log didn't often show up because we usually dealt with e as a base. And ln(e)=1 so it was ignored.

OpenStudy (richyw):

oh right. haha. now I feel dumb.

zepdrix (zepdrix):

And then you'd have to also expand out that log,\[\Large\bf\sf \log(1+\mathcal i)\quad=\quad \ln|1+\mathcal i|+\arg(1+\mathcal i)\]

OpenStudy (richyw):

yeah I could. but the answer you did is correct.

zepdrix (zepdrix):

And ummmm define a branch cut somewhere that works for the differentation or something? Blah I haven't looked at my notes enough yet..

zepdrix (zepdrix):

Oh ok :) heh

OpenStudy (richyw):

the question says assume appropriate branch cuts where neccesary, so I can just assume this is analytic. I just forgot my first year calc haha

zepdrix (zepdrix):

Ohhh interesting! :)

OpenStudy (richyw):

when the value of log(z) is specified, \(c^z\) is an entire function and \((c^z)' = c^z \log(z)\). It was sitting right there in my textbook! thanks a lot!

zepdrix (zepdrix):

Oh how bout that! np \c:/